### v1

@@ 4.2 Learning Convex Regularizers
where $C:=\{(p,q)|\psi_{w}(p)\leq q\}$, and $\delta_{C}$ denotes the indicator function of $C$. Here the linear operator $A$ denotes the X-ray transform with the prescribed geometry of the measurement setting and $u$ is the sinogram, representing the measured projections. For notational convenience, let $f_{2}(p,q)=\delta_{C}(p,q),f_{3}(v)=\gamma\psi_{w}(v)$. We then dualize the data fidelity in (32) which gives rise to the saddle-point problem:
@@
Figure 4 shows the qualitative results of reconstructing the test image in the sparse-view measurement scenario for all the above described methods as well as the corresponding ground truth image. Visually, the reconstruction results of the 2-layer ICNN trained using our proposed bilevel learning approach, as well as those from the learned primal-dual (LPD) method, exhibit high-quality reconstructions. They avoid both oversmoothed artifacts and overly fine-grained structures, which can often occur with other reconstruction methods. To support these results, we also show quantitative results in Table 1, which contains peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) for all methods on the employed test phantom. Note that this underlines the qualitative impressions from Figure 4. While the best scores are obtained using LPD, also the ICNN regularizer trained in our bilevel settings shows very competitive results. To some extent this can also be explained the number of learnable parameters which is in a ten-fold increase for LPD. On the other hand, the Fields of Experts regularizer that was trained using Algorithm 3 shows remarkable reconstruction results given its reduced complexity and limited number of learnable parameters. Even more interestingly, using the exact same architecture for the 2-layer ICNNs, our bilevel learning framework manages to outperform ACR by a large margin. Since they have identical learnable parameters, this shows very well the efficiency of the employed bilevel learning framework using primal-dual style differentiation. Overall the experiments show competitive results of FoE and ICNN-based regularizers trained using Algorithm 3. Similar results can also be expected in more challenging scenarios such as limited angle measurement settings, and the entire framework can be extended directly to other linear inverse problems with data-driven and potentially non-smooth regularizers.
@@ 5 Conclusions
For future work, extending our analysis to non-smooth variational problems would be a valuable direction. Furthermore, adapting this framework to stochastic settings could offer significant speed-ups and enable application to larger data sets. Finally, exploring the connections between this framework and other inexact methods, such as implicit function theorem-based approaches with linear solvers and automatic differentiation, represents another promising avenue for research.

### v2

@@ 4.2 Learning Convex Regularizers
where $C:=\{(p,q)|\psi_{w}(p)\leq q\}$, and $\delta_{C}$ denotes the indicator function of $C$. Here the linear operator $A$ denotes the X-ray transform with the prescribed geometry of the measurement setting and $u$ is the sinogram, representing the measured projections. To ensure strong convexity of the primal function $g$, we introduce a quadratic regularization term weighted by $\mu_{g}$. In practice, we select a small value (e.g., $\mu_{g}=10^{-8}$) to satisfy the strong convexity requirement without notably influencing the obtained results. For notational convenience, let $f_{2}(p,q)=\delta_{C}(p,q),f_{3}(v)=\gamma\psi_{w}(v)$. We then dualize the data fidelity in (33) which gives rise to the saddle-point problem:
@@
Figure 4 shows the qualitative results of reconstructing the test image in the sparse-view measurement scenario for all the above described methods as well as the corresponding ground truth image. Visually, the reconstruction results of the 2-layer ICNN trained using our proposed bilevel learning approach, as well as those from the learned primal-dual (LPD) method, exhibit high-quality reconstructions. They avoid both oversmoothed artifacts and overly fine-grained structures, which can often occur with other reconstruction methods. To support these results, we also show quantitative results in Table 1, which contains peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) for all methods on the employed test phantom. Note that this underlines the qualitative impressions from Figure 4. While the best scores are obtained using LPD, also the ICNN regularizer trained in our bilevel settings shows very competitive results. To some extent this can also be explained the number of learnable parameters which is in a ten-fold increase for LPD. On the other hand, the Fields of Experts regularizer that was trained using Algorithm 3 shows remarkable reconstruction results given its reduced complexity and limited number of learnable parameters. Even more interestingly, using the exact same architecture for the 2-layer ICNNs, our bilevel learning framework manages to outperform ACR by a large margin. Since they have identical learnable parameters, this shows very well the efficiency of the employed bilevel learning framework using primal-dual style differentiation. Overall the experiments show competitive results of FoE and ICNN-based regularizers trained using Algorithm 3. Future work could investigate the performance of the proposed method in more challenging scenarios, such as limited-angle measurement settings. Furthermore, the entire framework can be directly extended to other linear inverse problems with data-driven and potentially non-smooth regularizers.
@@ 5 Conclusions
For future work, extending our analysis to non-smooth variational problems would be a valuable direction. Furthermore, adapting this framework to stochastic settings could offer significant speed-ups and enable application to larger data sets. A comparison between our framework and single-level methods, particularly [32], which shares a similar type of lower-level structure, could be a promising direction for future work. This comparison could be further extended to fully first-order methods [33, 34, 35], providing deeper insights into their relative advantages and limitations. Finally, exploring the connections between this framework and other inexact methods, such as implicit function theorem-based approaches with linear solvers and automatic differentiation, represents another promising avenue for research.